I am having troubles with a couple problems:
solve the linear equation p/3-3p/8=3
I do not know why I can't get this. Thanks in advance!!

Answer by mananth(16949)

I am having troubles with a couple problems:
solve the linear equation p/3-3p/8=3
I do not know why I can't get this.
p/3 - 3p/8 =3
take 24 as LCD
8p-9p / 24 = 3
-p = 24*3
P=-72